What is the pH for a 0.30M weak base solution with Kb 3.7 x 10-6 1) 9.5x 10-12 2) 0.0011 3) 2.98 4) 9.5 5) 11.02
Answer #1

B dissociates as:

B +H2O -----> BH+ + OH-

0.3 0 0

0.3-x x x

Kb = [BH+][OH-]/[B]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((3.7*10^-6)*0.3) = 1.054*10^-3

since c is much greater than x, our assumption is correct

so, x = 1.054*10^-3 M

So, [OH-] = x = 1.054*10^-3 M

use:

pOH = -log [OH-]

= -log (1.054*10^-3)

= 2.9773

use:

PH = 14 - pOH

= 14 - 2.9773

= 11.02

Answer: 11.02